5. If 32.5 mL of 0.72 M NaOH is used to titrate a 0.20 M solution of HC,H302 what was the volume of HC2H3O2 used?
Answer #1

Concentration of NaOH solution = 0.72 M = 0.72 mol/L

Volume of NaOH solution = 32.5 ml = 32.5 L / 1000 = 0.0325 L

Number of moles of NaOH = 0.72 mol/L * 0.0325 L = 0.0234 mol

Concentration of HC2H3O2 solution = 0.20 M = 0.20 mol/L

Reaction:- NaOH(aq) + HC2H3O2 (aq) ---> NaC2H3O2(aq) + H2O (l)

1.0 mol of NaOH required 1.0 mol of HC2H3O2 then 0.0234 mol of NaOH will required 0.0234 mol of HC2H3O2.

Volume of HC2H3O2 solution required = 0.0234 mol / 0.20 mol/L = 0.117 L = 0.117 * 1000 ml = 117 ml

Volume of HC2H3O2 used = 117 ml